I need help with this question: You purchased a piece of property for $30,000 nine years ago and sold it today at $83,190. What is the annual rate of return on your investment? Show all work
Answer #1

Purchase price P = 30,000; Selling price S = 83,190; number of holding years n = 9

Total return = (S/P) -1 = (83,190/30,000) -1 = 177.30%

Annual return = [(1+ 177.30%)^(1/9)] -1 = 12.00%